Please help me solve. Factor completely. Thank you
y^4+3y^3-5y^2-15y

Answer by stanbon(75887)

y^4+3y^3-5y^2-15y
Factor out the "y" factor:
=y(y^3+3y^2-5y-15)
Factor the 1st two and the last two terms separately to get:
=y(y^2(y+3)-5(y+3))
Factor out the "y+3" factor
=y(y+3)(y^2-5)
===================
Cheers,
Stan H.